OpenStudy (anonymous):

What is the solution of the system? Use a graph. y = x + 5 y = –5x – 1

OpenStudy (anonymous):

Today is a really hard day for me

OpenStudy (anonymous):

(-1,4)

OpenStudy (anonymous):

Use substitution by substituting in the value for y. y = x + 5 -5x - 1 = x + 5 -6x = 6 x = -1 Then plug it in to get y. y = x + 5 y = -1 + 5 y = 4 (-1, 4)

OpenStudy (anonymous):

I can put the graphs on here

OpenStudy (anonymous):

Well that didn't work lol

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You see, you graphed it wrong. It says y = x + 5. You go up 5, not down 5.

OpenStudy (anonymous):

Yea I see that now I always tend to go down first when it comes to graphs I don't know why I do that